OpenStudy (anonymous):

how about this Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $25.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 100 1000 500 50 Answer: 50 .6x+25=.5x+30 -.5x -.5x ------------------------ .1x+25=30 -25 -25 ------------------ .1x=5 x=50 OR $0.05 x 500 = $25 + $30 = $55 $0.06 x 500 = $30 + $25 = $55 500 minutes. help

OpenStudy (mathstudent55):

6 cents is 0.06, not 0.6

OpenStudy (anonymous):

my bad lol

OpenStudy (mathstudent55):

You are on the right track, but be careful with little details. This is the equation you must solve. 0.06x+25=0.05x+30

OpenStudy (anonymous):

but i got to differ answers

OpenStudy (mathstudent55):

0.06x + 25 = 0.05x + 30 -0.05x =0.05x -------------------------- 0.01x + 25 = 30 -25 -25 -------------------------- 0.01x = 5 /0.01 /0.01 x = 500

OpenStudy (anonymous):

ok i get what you mean so far.

OpenStudy (anonymous):

my first answer was right not my 2nd.

OpenStudy (anonymous):

first one is not right but my 2nd one is

OpenStudy (mathstudent55):

Right, now you got it. All you need to do is solve the correct equation just like you solved your equation. You solved your equation correctly. Your mistake was simply using 0.6 and 0.5 instead of 0.06 and 0.05. Your second answer is correct. It's 500 minutes.

OpenStudy (anonymous):

ok i see now

OpenStudy (anonymous):

can you help me on this The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown.